Two ordinary fair dice are rolled. Work out the probability that both land on a value less than 4.

The probability of an event= Number of ways an event can occur/ Total number of possible outcomesBoth dices land on values less than 4. This means they both can be 1,2 or 3. Therefore, number of ways the even can occur is in this case 33=9. Total possible outcomes include all the combinations between 1,2,3,4,5 and 6 (the values on the first dice) and 1,2,3,4,5 and 6 (the values on the second dice). It means that the total number of possible outcomes is 66=36. The probability is thus 9/36=1/4=25%.

GS

Related Maths GCSE answers

All answers ▸

The equation of line L is y= 3x+2 and the equation of line M is 3y–9x+5=0. Show that these lines are parallel.


A sequence starts with the following terms... 2, 8, 14, 20... find the nth term


The equation of the line L1 is y=3x–2. The equation of the line L2 is 3y–9x+5=0. Show that these two lines are parallel.


Solve n^2+n-90=0 to find the value of n?